# Fast sampling and model selection for Bayesian mixture models

### v1

@@ A.4 Implementation
Table 3: The 32 questions selected from the CBS/New York Times poll for the data set used in this paper, listed in the order in which they were asked in the survey.
Appendix B Data sets
In this appendix we describe the data sets used in our examples.
B.1 CBS/New York Times poll
@@
B.2 Alzheimer’s diagnostic survey
This data set, which comes from Walsh Walsh06 , records the presence or absence of six symptoms of potential Alzheimer’s disease in 240 patients at the Mercer’s Institute national memory clinic in Dublin, Ireland, as reported by the patients’ primary caregivers on the occasion of the patients’ first visit to the clinic. The patients were preselected as having suspected Alzheimer’s disease or other age-related cognitive decline, but only mild symptoms, so that a positive diagnosis was not unlikely but also not guaranteed. The symptoms identified were: activity disturbance, affective disorder, aggression, agitation, diurnal rhythm disturbance, and hallucinations. The survey, a standard diagnostic instrument called the Behave-AD questionnaire, returns information on both the presence and the severity of each symptom, but Walsh converted the data to dichotomous presence/absence variables and this is the version we analyze. The data set is included in full in a table within the paper by Walsh.
@@ B.2 Alzheimer’s diagnostic survey
College education $<30\%$, 30–35%, 35–40%, $>40\%$
@@
Football Yes, no
Gun laws Yes, unlicensed, licensed, concealed only
Medicaid Yes, no
Sales tax Low, medium, high
Same-sex marriage Allowed, banned by statute
Temperature Low, medium, high
@@
Table 4: Summary of the 12 variables used in the 50 states data set. “College education” measures the fraction of the population over the age of 25 with a bachelor’s degree. “Football” refers to whether the state is home to a team in the US National Football League. “Gun laws” distinguishes constitutional carry, concealed carry plus unlicensed open carry, concealed carry plus licensed open carry, and concealed carry only. “Medicaid” refers to adoption of the Medicaid expansion under the Affordable Care Act of 2010, at the time of writing of this paper. “Sales tax” divides sales tax rates into three bands: low (below 1.75%), medium (1.75–7.07%), and high (above 7.07%). “Same-sex marriage” refers to state statutes and constitutions that contain, or do not contain, wording allowing or forbidding same-sex marriage. Actual same-sex marriage has been legal nationwide in the US since 2015 by federal law, which preempts state laws. “Temperature” records annual average temperatures in three bands: low (below $5^{\circ}$), medium (5–$15^{\circ}$), and high (above $15^{\circ}$).
B.3 Problem behaviors among teenagers
This data set comes from Li et al. LLSL18 and describes self-reported incidence of problem behaviors by 6504 teenagers in the United States during the 1990s. The behaviors probed were: lying to parents, unruly public behavior, damaging property, stealing from a store, stealing something worth less than US\$50, and taking part in a fight. Each behavior is recorded simply as present or absent, so there are only two possible responses for each survey item. The data set is included in full in a table within the paper by Li et al.
B.4 50 states data set
This data set, which was assembled by the current author from online sources for the purposes of this study, records 12 traits for each of the 50 United States (but excluding the District of Columbia, Puerto Rico, and other non-state territories). The 12 traits are listed in Table 4 and further details are given in the table caption. The full data set is available for download from https://umich.edu/~mejn/lca.
@@ B.3 Problem behaviors among teenagers
Table 5: Summary of the 14 variables in the online dating data set. Missing data for each question is also coded as an additional possible answer.
B.5 Online dating
This data set contains dating profiles for 59 946 users of the online dating service OKCupid, as posted at https://www.kaggle.com/datasets/subhamyadav580/ dating-site by Shubham Yadav. The full data set contains both textual descriptions contributed by the users and answers to multiple-choice profile questions, but we focus on the latter only, and on 14 questions in particular, which are listed in Table 5. We also include a “missing data” option as an additional possible response for each question, since significant numbers of people do not answer all questions.

### v2

@@ D.2 Candy dispenser
D.3 CBS/New York Times poll
@@
D.4 Alzheimer’s diagnostic survey
This data set, which comes from Walsh [60], records the presence or absence of six symptoms of potential Alzheimer’s disease in 240 patients at the Mercer’s Institute national memory clinic in Dublin, Ireland, as reported by the patients’ primary caregivers on the occasion of the patients’ first visit to the clinic. The patients were preselected as having suspected Alzheimer’s disease or other age-related cognitive decline, but only mild symptoms, so that a positive diagnosis was not unlikely but also not guaranteed. The symptoms identified were: activity disturbance, affective disorder, aggression, agitation, diurnal rhythm disturbance, and hallucinations. The survey, a standard diagnostic instrument called the Behave-AD questionnaire, returns information on both the presence and the severity of each symptom, but Walsh converted the data to dichotomous presence/absence variables and this is the version we analyze. The data set is included in full in a table within the paper by Walsh.
Table 4: Summary of the 12 variables used in the 50 states data set. “College education” measures the fraction of the population over the age of 25 with a bachelor’s degree. “Football” refers to whether the state is home to an American football team in the US National Football League. “Gun laws” distinguishes constitutional carry, concealed carry plus unlicensed open carry, concealed carry plus licensed open carry, and concealed carry only. “Medicaid” refers to adoption of the Medicaid expansion under the Affordable Care Act of 2010, at the time of writing of this paper. “Same-sex marriage” refers to state statutes and constitutions that contain, or do not contain, wording forbidding same-sex marriage. Actual same-sex marriage has been legal nationwide in the US since 2015 under federal law, which preempts state laws.
@@ D.4 Alzheimer’s diagnostic survey
College education Fraction with bachelor’s: below $30\%$, 30–35%, 35–40%, above $40\%$
@@
Football Has an NFL team, does not have an NFL team
Gun laws Yes, unlicensed open carry, licensed open carry, concealed carry only
Medicaid Expanded under the ACA, not expanded
Sales tax Below 1.75%, 1.75–7.07%, above 7.07%
Same-sex marriage Banned by statute or constitution, not banned
[…]
@@
D.5 50 states data set
This data set, which was assembled by the current author from online sources for the purposes of this study, records 12 traits for each of the 50 United States (but excluding the District of Columbia, Puerto Rico, and other non-state territories). The 12 traits are listed in Table 4 and further details are given in the table caption. The full data set is available for download from https://umich.edu/˜mejn/mixture.
D.6 Problem behaviors among teenagers
This data set comes from Li et al. [30] and describes self-reported incidence of problem behaviors by 6504 teenagers in the United States during the 1990s. The behaviors probed were: lying to parents, unruly public behavior, damaging property, stealing something worth less than US\$50, stealing from a store, and taking part in a fight. Each behavior is recorded simply as present or absent, so there are only two possible responses for each survey item. The data set is included in full in a table within the paper by Li et al.
Table 5: Summary of the 14 variables in the online dating data set. Missing data is also coded as an additional possible answer for each question.
@@ D.6 Problem behaviors among teenagers
D.7 Online dating
This data set contains dating profiles for 59 946 users of the online dating service OKCupid, as posted at https://www.kaggle.com/datasets/subhamyadav580/ dating-site by Shubham Yadav. The full data set contains both textual descriptions contributed by the users and answers to multiple-choice profile questions, but we focus on the latter only, and on 14 questions in particular, which are listed in Table 5. We include a “missing data” option as an additional possible response for each question, since significant numbers of people did not answer all questions.